OpenStudy (anonymous):

Help!!! Please!! Ill give out medals! Evaluate (f*h)(a) and (h*f)(1) f(x)=3x-1, g(x)=1/x+2, h(x)=square root of 4-x. Please help!

OpenStudy (anonymous):

@wio please help me if you can!

OpenStudy (anonymous):

function composition, right?

OpenStudy (anonymous):

First find \(f(a)\)

OpenStudy (anonymous):

Can you do that?

OpenStudy (anonymous):

\[ f(a) = 3a-1 \]

OpenStudy (anonymous):

ok im gona go step by step with you. Cause i get the concept but am really confused by this problem inparticular.

OpenStudy (anonymous):

\[ h(a) = \sqrt{4-a} \]\[ f(h(a)) = f(\sqrt{4-a}) = 3\sqrt{4-a}-1 \]

OpenStudy (anonymous):

Wait, are you evaluating for f*h. Im sorry. Not sure if you are but its f of h of a. and h of f of 1.

OpenStudy (anonymous):

I just did f of h of a

OpenStudy (anonymous):

\[ f(1) - 3(1)-1=2 \]\[ h(f(1)) = h(2) = \sqrt{4-2}=\sqrt 2 \]

OpenStudy (anonymous):

You can do it.

OpenStudy (anonymous):

thanks a lot man! Im trying. Just not having fun with this. Logic on zero sleep, lol.

OpenStudy (anonymous):

Hey ummm.. so in the end, i combine rad 2 and the end result of 3 rad 4-a-1?